# Thermodynamic properties of the macroscopically degenerate tetramer-dimer phase of the spin-1/2 Heisenberg model on the diamond-decorated square lattice

### v1

@@ Abstract
The spin-1/2 Heisenberg antiferromagnet on the diamond-decorated square lattice in the presence of a magnetic field displays various quantum phases including the Lieb-Mattis ferrimagnetic, dimer-tetramer, monomer-dimer, and spin-canted phases, in addition to the trivial fully saturated state. Thermodynamic properties of this model are investigated using several complementary analytical and numerical methods such as exact diagonalization up to the systems of 40 spins, an effective monomer-dimer description, sign-problem-free quantum Monte Carlo simulations for up to 180 spins, and a decoupling approximation. Our particular attention is focused on the parameter region favoring the dimer-tetramer phase. This ground state can be represented by a classical hard-dimer model on the square lattice and retains a macroscopic degeneracy even under a magnetic field. However, the description of the low-temperature thermodynamics close to the boundary between the macroscopically degenerate dimer-tetramer and the non-degenerate monomer-dimer phases requires an extended classical monomer-dimer lattice-gas model. Anomalous thermodynamic properties emerging in the vicinity of the dimer-tetramer phase are studied in detail. Under the adiabatic demagnetization we detect an enhanced magnetocaloric effect promoting an efficient cooling to absolute zero temperature, provided that the system reaches the dimer-tetramer ground state at zero field.
@@ I Introduction
Frustrated quantum spin systems constitute a challenge for theory, for example because the infamous sign problem [29, 30] precludes straightforward and efficient quantum Monte Carlo simulations. For this reason, tailor-made models such as the Rokhsar-Kivelson [31] and Kitaev [32] models have been proposed in order to allow for obtaining an exact solution. However, geometric frustration can sometimes also be beneficial rather than detrimental: if kinetic energy is significantly or even completely suppressed by destructive quantum interference, the quantum problem may be mapped to an effective classical dimer model [33] or at least well approximated by a quantum dimer model [34], thus rendering a controlled approximation if not an exact treatment of the low-energy physics possible. Examples of such phenomena include but are not limited to models with exact dimer ground states in one [16] and two dimensions [17], the one-third plateau in the kagome quantum antiferromagnet [35, 36], and localized magnons appearing just below the saturation field of highly frustrated antiferromagnets [37, 38, 39, 40, 41, 42].
Figure 1: The spin-1/2 Heisenberg model on the diamond-decorated square lattice with one unit cell shown by the yellow square. The decoupling approximation of Sec. III.2 is developed from a spin-star cluster framed in a blue square.
@@ II Heisenberg model on the diamond-decorated square lattice
In the following, we consider the spin-1/2 Heisenberg model on the diamond-decorated square lattice depicted in Fig. 1, and defined through the Hamiltionian
@@
The spin-1/2 operator $\bm{S}_{i,\mu}=({S}_{i,\mu}^{x},{S}_{i,\mu}^{y},{S}_{i,\mu}^{z})$ is assigned to the $\mu$th spin of the $i$th unit cell. The indices $i-\hat{x}$ and $i-\hat{y}$ refer to the unit cells immediately to the left and below the $i$th unit cell, respectively. We consider a finite lattice consisting of $N$ unit cells, containing $N_{s}=5N$ spins, and we apply periodic boundary conditions. Typically, we utilize square lattices with lattice sizes $L_{x}$, $L_{y}$, and the total number of unit cells $N=L_{x}L_{y}$. The two distinct exchange interactions $J_{1}$ and $J_{2}$ are depicted in Fig. 1 by thin and thick lines, respectively. The magnetic-field ($h$) term represents the standard Zeeman coupling of the spin degrees of freedom to an external magnetic field.
Figure 2: (a) Ground-state phase diagram of the spin-1/2 Heisenberg model on the diamond-decorated square lattice. The following phases were obtained in Ref. [48]: saturated paramagnetic (PM), spin-canted (SC), Lieb-Mattis (LM) ferrimagnetic, dimer-tetramer (DT), and monomer-dimer (MD) phases. Panel (b) zooms in on the DT phase that is the focus of the present work and provides cartoons of representative ground states in the DT, MD, and LM phases. Red broken lines show phase boundaries as obtained from the spin-star decoupling approximation discussed in Sec. III.2.
@@
where $|s\rangle_{d}=(|\!\!\uparrow\rangle|\!\!\downarrow\rangle-|\!\!\downarrow\rangle|\!\!\uparrow\rangle)/\sqrt{2}$ represents the dimer-singlet state. The DT phase is particularly intriguing because it comprises a macroscopic number of states that contain both singlet dimers and singlet tetramers. The singlet tetramers contain in turn triplet dimers, as depicted in Fig. 2 by small and large yellow ovals, respectively. Furthermore, each singlet tetramer is surrounded by singlet dimers within the DT ground-state manifold. As a result, overall these DT states are singlets and are of the following form:
@@ III.1 Effective monomer-dimer (EMD) model
Here we have introduced a distinction between horizontal ($H$) and vertical dimers ($V$) with two different interaction constants $J_{2}^{H}$ and $J_{2}^{V}$ in anticipation of a transfer-matrix approach that will treat the two spatial directions differently.
It should be noted that the two monomer spins in a diamond do not contribute to the Zeeman term in case they are included in the singlet tetramer. As a result, the effective low-temperature model can be considered as a monomer-dimer problem on the square lattice [49], where dimers can be located on the bonds of a square lattice connecting nearest-neighbor vertices (which corresponds to the singlet tetramer state of the diamond-decorated square lattice) and each vertex may host no more than one dimer (i.e., different singlet tetramers cannot have common spins). This correspondence is illustrated in Figs. 3(a), (b). Defining the energy with respect to the dimer singlets, the partition function of such a model can be written as follows:
@@
Figure 3: Schematic of a particular configuration of the mixture of MD and DT phases in different representations: (a) the diamond-decorated square lattice where yellow ovals indicate a singlet state on the highlighted clusters; (b) EMD representation of the same model; (c) the spin representation for the transfer-matrix formalism [49].
In order to calculate the partition function, we resort to the transfer-matrix formalism [49] (see also Refs. [54, 61, 55]). The transfer matrix proceeds on two neighboring rows of vertical bonds. Vertical dimers are identified with spins up ($\uparrow$), and the absence of dimers on vertical bonds corresponds to spin down ($\downarrow$) (see Fig. 3(c)). The transfer matrix in operator form contains three contributions, ${\cal V}=V_{3}V_{2}V_{1}$, with
@@ III.2 Spin-star decoupling approximation
Figure 4: A spin-star fragment of the spin-1/2 Heisenberg model on the diamond-decorated square lattice (left), which can be mapped to a simpler composite spin-star due to local conservation of the total spin on each dimer (right).
The spin-star decoupling approximation is developed from the idea of decomposing the total Hamiltonian of the spin-1/2 Heisenberg model on the diamond-decorated square lattice (1) into a sum of non-commuting cluster Hamiltonians ${\cal H}^{0}=\sum_{j=1}^{N}{\cal H}_{j}^{0}$ ascribed to individual spin-star clusters. As a starting point of our calculation we will therefore consider the Hamiltonian ${\cal H}_{j}^{0}$ of a single spin-star cluster, which is composed of one monomeric spin and its four neighboring dimers. The considered spin-star cluster is schematically illustrated on the left-hand-side of Fig. 4 including simplified notation and is mathematically defined as follows
@@
The inclusion of the factor of 1/2 in the second term ensures that the intra-dimer interaction $J_{2}$ will not be double-counted if one performs a summation over the spin-star Hamiltonians $H_{j}^{0}$ at the end of our calculation. Note furthermore that we have omitted the Zeeman term because this term can be trivially added at the end of the calculation as it only shifts the respective eigenenergies thanks to the total Hamiltonian and the Zeeman term commuting with each other. Within the spin-star cluster decoupling, the total energy of the spin-1/2 Heisenberg diamond-decorated square lattice will be obtained as the sum of eigenenergies of the spin stars given by the cluster Hamiltonian (17) upon simply disregarding the non-commutative nature of the cluster Hamiltonians $[{\cal H}_{i}^{0},{\cal H}_{j}^{0}]\neq 0$ for neighbors $i$ and $j$.
@@
which can be further simplified into the following compact form
@@
Summing the eigenenergies of the spin stars and adding the respective Zeeman term yields the following formula for the overall energy of the spin-1/2 Heisenberg model on the diamond-decorated square lattice $E_{T}(\{n_{i}\})=E_{T}^{0}(\{n_{i}\})-hS_{T}^{z}$, where $E_{T}^{0}(\{n_{i}\})$ represents the respective zero-field energy depending on the set of all occupation numbers $\{n_{i}\}$:
@@
The free energy of the spin-1/2 Heisenberg model on the diamond-decorated square lattice within the spin-star decoupling approximation can then be calculated according to the formula:
@@
The spin-star decoupling approximation recovers the transition fields around the DT phase in the ground-state phase diagram (see red broken lines in Fig. 2(b)). By setting all $n_{i}=0$ in Eq. (22), we actually get the exact value for the ground-state energy of the MD phase (8). Requiring that there is a single triplet dimer in each spin-star configuration, we instead obtain the exact energy of the DT phase (9). Calculating the ground-state energy of the LM phase (all $n_{i}=1$) within the spin-star approach, $E_{\rm LM}=-\frac{N}{2}(5J_{1}-J_{2}+3h)$, yields the transition field between the LM and DT phases as $h=J_{2}-J_{1}$, which is very close to the DMRG results shown in Fig. 2(b) (cf. the red broken and blue solid lines in Fig. 2(b)). Recall that on the phase boundary the energy of any configuration that contains at least one triplet in the spin-star cluster has the same energy. This leads to an additional macroscopic degeneracy (higher than the degeneracy of the DT phase). This artifact of the decoupling approximation influences the low-temperature behavior entropy discussed in Sec. IV.
The decoupling approximation based on the spin-star clusters was developed under two fundamental assumptions, (i) ignoring the non-commutative nature of the cluster Hamiltonians (17) and (ii) subsequently considering only the five energetically most favorable states of the spin star given by Eq. (III.2). Despite these simplifications, the spin-star cluster decoupling offers a reasonable approximation at sufficiently low temperatures whenever the system is driven towards the fragmented MD or DT phases, whereas the collective nature of the LM ferrimagnetic phase is not rigorously accounted for within this treatment. Although we have attempted to improve precision of the spin-star approximation in zero magnetic field by considering degeneracies pertinent to free monomeric spins, this calculation procedure should be regarded as much more reliable and precise in the presence of a magnetic field due to Zeeman splitting of the relevant spin multiplets. It indeed turns out that the quantitative agreement between the results derived from the spin-star approximation and a numerical treatment of the full model progressively extends to higher temperatures as the magnetic field increases, see Sec. IV.
@@ III.3 Exact diagonalization
The full $N_{s}=30$ spectra were obtained previously [48] and can be re-evaluated for the present purposes. However, here we push exact diagonalization (ED) further to the $N_{s}=40$ system sketched in Fig. 6. We have followed the same strategy as in Ref. [48], exploited conservation of the total spin on each dimer and thus passed to an effective Lieb lattice, represented by the yellow spheres in the background of Fig. 6. For the present $N_{s}=40$ system, computer enumeration then yields 191 topologically nonequivalent arrangements of spin triplets and singlets on the dimers with degeneracies ranging from 1 (for all dimers in the triplet state) to 1152. We were then able to compute the full spectra in the sectors with up to $N_{\rm trip.}=8$ triplets. This includes in particular the sector with $N_{\rm trip.}=4$ triplets that contains the dimer-tetramer phase for the $N_{s}=40$ geometry. For the sectors with 9 triplets or more, we further diagonalized high-$S_{z}$ sectors completely and obtained low-energy levels using the Lanczos procedure [62] combined with the strategy of Ref. [63]. Specifically, in the Lieb-Mattis regime that corresponds to 16 triplets in the $N_{s}=40$ system, we were only able to obtain full spectra for $S^{z}\geq 13$ and had to contend ourselves with low-lying levels for $S^{z}\leq 12$, i.e., the ferrimagnetic state with magnetization $3/5$ and below. Furthermore, we reconstructed the spin multiplets using $S^{z}$ classification and spin inversion symmetry in the $S^{z}=0$ sector; the latter permits us to avoid diagonalizations in the $S^{z}=1$ sector, which would usually have the highest dimension. All in all, we were able to recover an entropy of $36.082\,\ln(2)$, i.e., 90.2% of the total entropy of the $N_{s}=40$ system.
@@
Figure 7 shows the gaps to the lowest excitation above the DT ground-state for the $N_{s}=40$ system ($N_{\rm trip.}=4$). For $J_{2}\gtrsim 1.5J_{1}$, the gap to the $N_{\rm trip.}=3$ excitation coincides with the DT-MD phase transition in the phase diagram Fig. 2, as in fact in this case states with any number $N_{\rm trip.}\leq N_{s}/10$ collapse. On the other hand, the field-induced DT-LM transition in Fig. 2 is a true first-order transition. Thus, the closing of the $N_{\rm trip.}=5$ gap in Fig. 7 is preempted by a direct transition to a $N_{\rm trip.}=16$ configuration, reproduced from Fig. 2 by the dashed line in Fig. 7. As a consequence, the maximum of the gap is found for $J_{2}/J_{1}\approx 1.4$ whereas the maximal extension of the DT phase in a magnetic field occurs for $J_{2}/J_{1}\approx 1.5$. Furthermore, one observes an $N_{\rm trip.}=6$ bound state coming down in energy when approaching the DT-LM transition at $h=0$. The nature of the DT-LM transition at $h=0$ is less clear to decide, at least based on the $N_{s}=40$ ED data, i.e., one does observe higher-$N_{\rm trip.}$ / higher-$S$ states coming down in energy, but it is not clear if they actually collapse in the thermodynamic limit, or if the DT-LM transition remains a conventional first-order transition at $h=0$.

### v2

@@ Abstract
The spin-1/2 Heisenberg antiferromagnet on the diamond-decorated square lattice in the presence of a magnetic field displays various quantum phases including the Lieb-Mattis ferrimagnetic, dimer-tetramer, monomer-dimer, and spin-canted phases, in addition to the trivial fully saturated state. Thermodynamic properties of this model are investigated using several complementary analytical and numerical methods such as exact diagonalization up to the systems of 40 spins, an effective monomer-dimer description, sign-problem-free quantum Monte Carlo simulations for up to 180 spins, and a decoupling approximation. Our particular attention is focused on the parameter region favoring the dimer-tetramer phase. This ground state can be represented by a classical hard-dimer model on the square lattice and retains a macroscopic degeneracy even under a magnetic field. However, the description of the low-temperature thermodynamics close to the boundary between the macroscopically degenerate dimer-tetramer and the non-degenerate monomer-dimer phases requires an extended classical monomer-dimer lattice-gas model. Anomalous thermodynamic properties emerging in the vicinity of the dimer-tetramer phase are studied in detail. Under the adiabatic demagnetization we detect an enhanced magnetocaloric effect promoting an efficient cooling to absolute zero temperature.
@@ I Introduction
Frustrated quantum spin systems constitute a challenge for theory—for example, because the infamous sign problem [29, 30] precludes straightforward and efficient quantum Monte Carlo simulations. For this reason, tailor-made models such as the Rokhsar-Kivelson [31] and Kitaev [32] models have been proposed in order to allow for obtaining an exact solution. However, geometric frustration can sometimes also be beneficial rather than detrimental: if kinetic energy is significantly or even completely suppressed by destructive quantum interference, the quantum problem may be mapped to an effective classical dimer model [33] or at least well approximated by a quantum dimer model [34], thus rendering a controlled approximation if not an exact treatment of the low-energy physics possible. Examples of such phenomena include but are not limited to models with exact dimer ground states in one [16] and two dimensions [17], the one-third plateau in the kagome quantum antiferromagnet [35, 36], and localized magnons appearing just below the saturation field of highly frustrated antiferromagnets [37, 38, 39, 40, 41, 42].
Figure 1: The spin-1/2 Heisenberg model on the diamond-decorated square lattice with one unit cell shown by the yellow square. The decoupling approximation of Sec. III.2 is developed from the spin-star cluster framed in a blue square.
@@ II Heisenberg model on the diamond-decorated square lattice
In the following, we consider the spin-1/2 Heisenberg model on the diamond-decorated square lattice depicted in Fig. 1, and defined through the Hamiltonian
@@
The spin-1/2 operator $\bm{S}_{i,\mu}=({S}_{i,\mu}^{x},{S}_{i,\mu}^{y},{S}_{i,\mu}^{z})$ is assigned to the $\mu$th spin of the $i$th unit cell. The indices $i-\hat{x}$ and $i-\hat{y}$ refer to the unit cells immediately to the left and below the $i$th unit cell, respectively. We consider a finite lattice consisting of $N$ unit cells, containing $N_{s}=5N$ spins, and we apply periodic boundary conditions. Typically, we utilize square lattices with lattice sizes $L_{x}$, $L_{y}$, and the total number of unit cells $N=L_{x}L_{y}$. The two distinct exchange interactions $J_{1}$ and $J_{2}$ are depicted in Fig. 1 by thin and thick lines, respectively. The magnetic field ($h$) term represents the standard Zeeman coupling of the spin degrees of freedom to an external magnetic field.
Figure 2: (a) Ground-state phase diagram of the spin-1/2 Heisenberg model on the diamond-decorated square lattice. The following phases were obtained in Ref. [48]: saturated paramagnetic (PM), spin-canted (SC), Lieb-Mattis (LM) ferrimagnetic, dimer-tetramer (DT), and monomer-dimer (MD) phases. Panel (b) zooms in on the DT phase that is the focus of the present paper and provides cartoons of representative ground states in the DT, MD, and LM phases. Red-broken lines show phase boundaries as obtained from the spin-star decoupling approximation discussed in Sec. III.2.
@@
where $|s\rangle_{d}=(|\!\!\uparrow\rangle|\!\!\downarrow\rangle-|\!\!\downarrow\rangle|\!\!\uparrow\rangle)/\sqrt{2}$ represents the dimer-singlet state. The DT phase is particularly intriguing because it comprises a macroscopic number of states that contain both singlet dimers and singlet tetramers. The singlet tetramers contain in turn triplet dimers, as depicted in Fig. 2 by large yellow and small purple ovals, respectively. Furthermore, each singlet tetramer is surrounded by singlet dimers within the DT ground-state manifold. As a result, overall these DT states are singlets and are of the following form:
@@ III Methods
At the beginning of each of the following subsections, we briefly outline its purpose. Readers primarily interested in the results and their implications may skip the technical details in the corresponding subsections, or proceed directly to Sec. IV, where our findings are discussed in detail.
@@ III.1 Effective monomer-dimer (EMD) model
In this subsection, we derive and discuss the transfer-matrix solution of the effective monomer-dimer (EMD) model as a simplified approach to describe the low-temperature thermodynamic properties of the spin-1/2 Heisenberg model in the DT phase. This model focuses on capturing the contributions of low-energy excitations from triplets to singlets and their influence on quantities such as entropy, specific heat, and magnetization.
@@
Here, we have introduced a distinction between horizontal ($H$) and vertical dimers ($V$) with two different interaction constants $J_{2}^{H}$ and $J_{2}^{V}$ in anticipation of a transfer-matrix approach that will treat the two spatial directions differently.
It should be noted that the two monomer spins in a diamond do not contribute to the Zeeman term in case they are included in the singlet tetramer. As a result, the effective low-temperature model can be considered as a monomer-dimer problem on the square lattice [49], where dimers can be located on the bonds of a square lattice connecting nearest-neighbor vertices (which corresponds to the singlet tetramer state of the diamond-decorated square lattice) and each vertex may host no more than one dimer ( i.e., different singlet tetramers cannot have common spins). This correspondence is illustrated in Figs. 3(a) and 3(b). Defining the energy with respect to the dimer singlets, the partition function of such a model can be written as follows:
@@
Figure 3: Schematic of a particular configuration of the mixture of MD and DT phases in different representations. (a) The diamond-decorated square lattice where yellow ovals indicate a singlet state on the highlighted clusters. (b) EMD representation of the same model. (c) The spin representation for the transfer-matrix formalism [49].
In order to calculate the partition function, we resort to the transfer-matrix formalism [49] (see also Refs. [54, 61, 55]). The transfer matrix proceeds on two neighboring rows of vertical bonds. Vertical dimers are identified with spins up ($\uparrow$), and the absence of dimers on vertical bonds corresponds to spin down ($\downarrow$) [see Fig. 3(c)]. The transfer matrix in operator form contains three contributions, ${\cal V}=V_{3}V_{2}V_{1}$, with
@@ III.2 Spin-star decoupling approximation
Figure 4: A spin-star fragment of the spin-1/2 Heisenberg model on the diamond-decorated square lattice (left), which can be mapped to a simpler composite spin-star owing to local conservation of the total spin on each dimer (right).
This subsection introduces the spin-star decoupling approximation as a tool to describe low-energy excitations and their thermodynamic effects from the short-range perspective. In contrast to the EMD approach of the previous subsection, here we also take triplet excitations on the dimers into account.
The spin-star decoupling approximation is developed from the idea of decomposing the total Hamiltonian of the spin-1/2 Heisenberg model on the diamond-decorated square lattice (1) into a sum of non-commuting cluster Hamiltonians ${\cal H}^{0}=\sum_{j=1}^{N}{\cal H}_{j}^{0}$ ascribed to individual spin-star clusters. As a starting point of our calculation we will therefore consider the Hamiltonian ${\cal H}_{j}^{0}$ of a single spin-star cluster, which is composed of one monomeric spin and its four neighboring dimers. The considered spin-star cluster is schematically illustrated on the left-hand side of Fig. 4 including simplified notation and is mathematically defined as follows:
@@
The inclusion of the factor of 1/2 in the second term ensures that the intradimer interaction $J_{2}$ will not be double-counted if one performs a summation over the spin-star Hamiltonians $H_{j}^{0}$ at the end of our calculation. Note furthermore that we have omitted the Zeeman term because this term can be trivially added at the end of the calculation as it only shifts the respective eigenenergies thanks to the total Hamiltonian and the Zeeman term commuting with each other. Within the spin-star cluster decoupling, the total energy of the spin-1/2 Heisenberg diamond-decorated square lattice will be obtained as the sum of eigenenergies of the spin stars given by the cluster Hamiltonian (17) upon simply disregarding the non-commutative nature of the cluster Hamiltonians $[{\cal H}_{i}^{0},{\cal H}_{j}^{0}]\neq 0$ for neighbors $i$ and $j$.
@@
which can be further simplified into the following compact form:
@@
Summing the eigenenergies of the spin stars and adding the respective Zeeman term yields the following formula for the overall energy of the spin-1/2 Heisenberg model on the diamond-decorated square lattice $E_{T}(\{n_{i}\})=E_{T}^{0}(\{n_{i}\})-hS_{T}^{z}$, where $E_{T}^{0}(\{n_{i}\})$ represents the respective zero-field energy depending on the set of all occupation numbers $\{n_{i}\}$,
@@
The free energy of the spin-1/2 Heisenberg model on the diamond-decorated square lattice within the spin-star decoupling approximation can then be calculated according to the formula
@@
The spin-star decoupling approximation recovers the transition fields around the DT phase in the ground-state phase diagram [see red broken lines in Fig. 2(b)]. By setting all $n_{i}=0$ in Eq. (22), we actually get the exact value for the ground-state energy of the MD phase (8). Requiring that there is a single triplet dimer in each spin-star configuration, we instead obtain the exact energy of the DT phase (9). Calculating the ground-state energy of the LM phase (all $n_{i}=1$) within the spin-star approach, $E_{\rm LM}=-\frac{N}{2}(5J_{1}-J_{2}+3h)$, yields the transition field between the LM and DT phases as $h=J_{2}-J_{1}$, which is very close to the DMRG results shown in Fig. 2(b) [cf. the red-broken and blue-solid lines in Fig. 2(b)]. Recall that on the phase boundary the energy of any configuration that contains at least one triplet in the spin-star cluster has the same energy. This leads to an additional macroscopic degeneracy (higher than the degeneracy of the DT phase). This artifact of the decoupling approximation influences the low-temperature behavior of the entropy discussed in Sec. IV.
The decoupling approximation based on the spin-star clusters was developed under two fundamental assumptions, (i) ignoring the non-commutative nature of the cluster Hamiltonians (17) and (ii) subsequently considering only the five energetically most favorable states of the spin star given by Eq. (III.2). Despite these simplifications, the spin-star cluster decoupling offers a reasonable approximation at sufficiently low temperatures whenever the system is driven towards the fragmented MD or DT phases, whereas the collective nature of the LM ferrimagnetic phase is not rigorously accounted for within this treatment. Although we have attempted to improve precision of the spin-star approximation in zero magnetic field by considering degeneracies pertinent to free monomeric spins, this calculation procedure should be regarded as much more reliable and precise in the presence of a magnetic field because of Zeeman splitting of the relevant spin multiplets. It indeed turns out that the quantitative agreement between the results derived from the spin-star approximation and a numerical treatment of the full model progressively extends to higher temperatures as the magnetic field increases, see Sec. IV.
@@ III.3 Exact diagonalization
Here, we provide some details on the exact diagonalization (ED) that we used to obtain numerically precise results for small system sizes. This method allows us to calculate a broad range of thermodynamic quantities, including magnetization, specific heat, and entropy, serving as a benchmark for validating analytical and approximate models.
The full $N_{s}=30$ spectra were obtained previously [48] and can be re-evaluated for the present purposes. However, here we push exact diagonalization (ED) further to the $N_{s}=40$ system sketched in Fig. 6. We have followed the same strategy as in Ref. [48], exploited conservation of the total spin on each dimer and thus passed to an effective Lieb lattice, represented by the yellow spheres in the background of Fig. 6. For the present $N_{s}=40$ system, computer enumeration then yields 191 topologically nonequivalent arrangements of spin triplets and singlets on the dimers with degeneracies ranging from 1 (e.g., for all dimers in the triplet state) to 1152. We were then able to compute the full spectra in the sectors with up to $N_{\rm trip.}=8$ triplets. This includes in particular the sector with $N_{\rm trip.}=4$ triplets that contains the dimer-tetramer phase for the $N_{s}=40$ geometry. For the sectors with nine triplets or more, we further diagonalized high-$S_{z}$ sectors completely and obtained low-energy levels using the Lanczos procedure [62] combined with the strategy of Ref. [63]. Specifically, in the Lieb-Mattis regime that corresponds to 16 triplets in the $N_{s}=40$ system, we were only able to obtain full spectra for $S^{z}\geq 13$ and had to contend ourselves with low-lying levels for $S^{z}\leq 12$, i.e., the ferrimagnetic state with magnetization $3/5$ and below. Furthermore, we reconstructed the spin multiplets using $S^{z}$ classification and spin inversion symmetry in the $S^{z}=0$ sector; the latter permits us to avoid diagonalizations in the $S^{z}=1$ sector, which would usually have the highest dimension. All in all, we were able to recover an entropy of $36.082\,\ln(2)$, i.e., 90.2% of the total entropy of the $N_{s}=40$ system.
@@
Figure 7 shows the gaps to the lowest excitation above the DT ground state for the $N_{s}=40$ system ($N_{\rm trip.}=4$). For $J_{2}\gtrsim 1.5J_{1}$, the gap to the $N_{\rm trip.}=3$ excitation coincides with the DT-MD phase transition in the phase diagram Fig. 2, as in fact in this case states with any number $N_{\rm trip.}\leq N_{s}/10$ collapse. On the other hand, the field-induced DT-LM transition in Fig. 2 is a true first-order transition. Thus, the closing of the $N_{\rm trip.}=5$ gap in Fig. 7 is preempted by a direct transition to a $N_{\rm trip.}=16$ configuration, reproduced from Fig. 2 by the dashed line in Fig. 7. As a consequence, the maximum of the gap is found for $J_{2}/J_{1}\approx 1.4$ whereas the maximal extension of the DT phase in a magnetic field occurs for $J_{2}/J_{1}\approx 1.5$. Furthermore, one observes an $N_{\rm trip.}=6$ bound state coming down in energy when approaching the DT-LM transition at $h=0$. The nature of the DT-LM transition at $h=0$ is less clear to decide, at least based on the $N_{s}=40$ ED data, i.e., one does observe higher-$N_{\rm trip.}$/higher-$S$ states coming down in energy, but it is not clear if they actually collapse in the thermodynamic limit, or if the DT-LM transition remains a conventional first-order transition at $h=0$.